# A Differentiable Sparse Vector Quantization (SVQ) for Spatio-Temporal Forecasting

### v3

@@
Abstract
Spatio-temporal forecasting, pivotal in numerous fields, hinges on the delicate equilibrium between isolating nuanced patterns and sifting out noise. To tackle this, we introduce Sparse Regression-based Vector Quantization (SVQ), a novel technique that leverages sparse regression for succinct representation, an approach theoretically and practically favored over classical clustering-based vector quantization methods. This approach preserves critical details from the original vectors using a regression model while filtering out noise via sparse design. Moreover, we approximate the sparse regression process using a blend of a two-layer MLP and an extensive codebook. This approach not only substantially cuts down on computational costs but also grants SVQ differentiability and training simplicity, resulting in a notable enhancement of performance. Our empirical studies on five spatial-temporal benchmark datasets demonstrate that SVQ achieves state-of-the-art results. Specifically, on the WeatherBench-S temperature dataset, SVQ improves the top baseline by 7.9%. In video prediction benchmarks—Human3.6M, KTH, and KittiCaltech—it reduces MAE by an average of 9.4% and improves image quality by 17.3% (LPIPS). Code is publicly available at https://anonymous.4open.science/r/SVQ-Forecasting.
1 Introduction
Spatio-temporal forecasting (Wang et al., 2022a) has board applications in many domains including environment, climate, transportation, etc. Many traditional methods relying on feature engineering have been studied in the past several decades. Recently, deep learning models are becoming more popular (Wang et al., 2022a). However, as more applications (e.g., weather forecasting) arise, it calls for deep models to accurately predict future dynamics with precise spatial and temporal resolutions.
Despite the various forms of spatio-temporal forecasting approaches, such as the use of graph neural network (GNN) models for unevenly resolved data, our work concentrates on grid-based datasets, one of the most widely adopted structures in the field. Within this context, we address a primary challenge: how to finely balance the preservation of important details against the need to filter out irrelevant noise, thereby enhancing the generalized performance of forecasting models. Although vector quantization (VQ) is effective in reducing noises by constraining latent patterns to a clustered codebook, it is usually limited in preserving the diversity of input patterns, making it suboptimal for spatial-temporal forecasting. We performed a comprehensive suite of experiments, the results of which indicate that, contrary to expectations, nearly all sota VQ methods we examined did not enhance the accuracy of a basic weather forecasting model as shown in Table 4. This outcome may provide insight into the infrequent application of VQ techniques in spatial-temporal prediction tasks.
To address this limitation of VQ, we propose a Sparse regression-based Vector Quantization (SVQ) method. By associating each vector to multiple codes from a large codebook, it is able to cover diverse visual patterns; by introducing a sparse regularization mechanism in the regression model, we effectively filter out noise and thus ensure a good generalized performance. As shown in Figure 1, our proposed SVQ method exhibits a more uniform distribution of quantized vectors compared to VQ, indicating that SVQ is able to preserve more diverse and fine-grained information from the original inputs than VQ.
Despite employing a series of vectors for representation instead of a single codebook vector, regression-based vector quantization remains true to the foundational aspects of the technique. It still quantizes input vectors into a finite set of discrete levels, utilizing a dynamic codebook-like structure. The key process of mapping input data to representative vectors is maintained, but with the added sophistication of regression models that adaptively combine vectors for a more accurate representation. This method preserves the essence of vector quantization—efficiently compressing and approximating data—while enhancing its ability to capture complex patterns within the input space.
To address the high computational cost of sparse regression, which scales quadratically with codebook size, we approximate the sparse regression model with a MLP model and a large codebook. It not only dramatically simplifies the computation, but more importantly, makes SVQ differentiable and thus easy to train. We emphasize the importance of differentiability of SVQ for its success in spatial-temporal forecasting. Our experiments reveal that lookup-based VQ methods can substantially increase prediction error, despite some employing a residual design that theoretically should perform on par with or better than non-VQ cases. This suggests the “straight-through” estimator, a non-differentiable component of VQ, may be a major factor behind VQ’s underperformance. In particular, the non-differentiable operator such as argmin or lookup introduces errors into gradient estimation of all modules preceding the quantization module, consequently undermining prediction accuracy. In contrast, our proposed SVQ is differentiable and easy to compute accurate gradients for training. We finally note that SVQ also differs from traditional VQ methods by not relying on intricate models to construct the prior distribution of codes, such as the GPT-style model in VQGAN (Esser et al., 2021) or the BERT-style model in MaskGIT (Chang et al., 2022).
Besides differentiability, there are two interesting properties of the proposed SVQ scheme. First, according to our empirical study, we find that with large enough codebook size, we can freeze the randomly initialized codebook without sacrificing any performance, making SVQ potentially more efficient for training. Secondly, unlike lookup-based methods, SVQ allows for both positive and negative weights, enhancing the ability to add or subtract patterns to improve representational power.
Finally, despite its simplification, our experiments across multiple real-world benchmark datasets—including weather forecasting, traffic flow prediction, and video forecasting, demonstrate the superior performance of our SVQ method. In WeatherBench-S temperature forecasting benchmark, it improves the prior SOTA method by 7.9%. For video prediction across renowned datasets—Human3.6M, KTH, and KittiCaltech—SVQ not only achieves a consistent 9.4% reduction in MAE but also enhances perceptual quality, evidenced by a 17.3% lower LPIPS score. These results highlight SVQ’s versatility and superior accuracy in diverse spatial-temporal prediction scenarios.
2 Related Work
2.1 Recurrent-based forecasting model
The majority of spatio-temporal forecasting models leverage techniques such as Conv2D (Xu et al., 2018), Conv3D (Wang et al., 2019a), and attention (Liu et al., 2022b) for spatial modeling. Distinctions among these models primarily arise from how they incorporate temporal information. Recurrent-based models, exemplified by ConvLSTM (SHI et al., 2015), have been widely used to capture motion dynamics by iteratively processing multi-frame predictions. Variants like PredRNN (Wang et al., 2017) introduce the Spatio-Temporal LSTM (ST-LSTM) unit, integrating spatial appearances and temporal variations within a single memory pool. Further advancements include PredRNN++ (Wang et al., 2018) and PredRNNV2 (Wang et al., 2022b), which deepen the model and expand the receptive field through a cascading LSTM mechanism and a memory decoupling strategy, respectively.
2.2 Non-recurrent forecasting model
Despite the effectiveness of recurrent-based methods, they suffer from high computational cost caused by their inherent unparallelizable architecture. Recent efforts in spatio-temporal forecasting have shifted towards non-recurrent models that decouple the forecasting task from autoregressive processes. Notably, SimVPv1 (Gao et al., 2022) and SimVPv2 (Tan et al., 2022) separate spatial and temporal learning into distinct phases within an encoder-translator-decoder structure, consistently surpassing recurrent counterparts in video prediction tasks. Due to their similar architectures, we will collectively refer to them as the SimVP model in the remainder of our paper. TAU (Tan et al., 2023a) refines the architecture by incorporating a visual attention mechanism into the translator. OpenSTL (Tan et al., 2023b) further enhances the translator by MetaFormers.
2.3 Vector quantization and sparse coding
@@
Instead of using continuous latent, VQVAE (van den Oord et al., 2017), a seminal work, integrates vector quantization (VQ) in learning discrete latent representations, typically assigning each vector to the nearest code in a codebook. Subsequent enhancements include Residual VQ (Zeghidour et al., 2022), which quantizes the residuals recursively, and Multi-headed VQ (Mama et al., 2021b), which adopts multiple heads for each vector. While effective, these methods often rely on a relatively small number of codes for representing continuous vectors. Addressing this, SCVAE (Xiao et al., 2023) employs sparse coding, allowing vectors to be represented through sparse linear combinations of multiple codes and achieving end-to-end training via the Learnable Iterative Shrinkage Thresholding Algorithm (LISTA) (Gregor & LeCun, 2010). However, a significant drawback of the sparse coding method using LISTA (dubbed SVQ-raw here) is its high computational complexity, which scales quadratically with codebook size.
@@
Our work is closely related to a simultaneous research work (Tschannen et al., 2023), which employs an infinite cookbook with a linear layer for continuous vector quantization applied in image generation. Despite the similarity, our approach is largely inspired by sparse regression, which is clearly evidenced by our analysis. In particular, we provide a better theoretical foundation and address the challenges arising from spatio-temporal forecasting well.
3 Sparse Vector Quantization (SVQ)
In this section, we will first describe the mathematical foundation for sparse vector quantization, followed by the description of detailed implementation.
3.1 Vector quantization by sparse regression
@@
where is the weight for combining codes to approximate . By introducing regularizer in the optimization problem, we effectively enforce to be associated with a small number of codes. Compared to classic VQ methods where codes have to be learned through clustering, according to (Chiu et al., 2022), it is sufficient to use randomly sampled vectors as codes as long as its number is large enough, thus avoiding the need of computing and adjusting codes. The theoretical guarantee of sparse regression is closely related to the property of subspace clustering, as revealed in Theorem 4.1.
@@
where , is the th element of vector , and outputs if and otherwise. We consider the first step of the iteration where and have and the resulting output for is given as
@@
3.2 Detailed implementation
Architecture of backbone model. As depicted in Figure 3, we adopt SimVP as our backbone model, which encompasses an encoder for spatial feature extraction, a translator for temporal dependency learning, and a decoder for frame reconstruction. We integrate the quantization module between the encoder and translator. The input data is a 4D tensor , representing height (), width (), time steps (), and channels (). The encoder condenses into downsampled latent representation , maintaining temporal dimensionality while altering spatial and channel dimensions. This latent space, composed of tokens, each a -dimensional vector, undergoes vector quantization.
Quantization module. SVQ comprises a two-layer MLP and a huge codebook. The codebook is a randomly initialized matrix , where denotes the size of codebook. To achieve automatic selection of codes, we generate a weight matrix by nonlinear projection from latent representation . This projection is formally expressed as , wherein the comprises two linear layers and an intermediate ReLU activation function. Subsequently, the quantized output is obtained by computing the dot product of weight matrix and codebook matrix , a process that can be conceptualized as a selection operation as shown in Figure 3. To encourage sparsity within the generated weight matrix, we apply a Mean Absolute Error (MAE) loss to the output as a surrogate form of regularization.
Despite the potential feasibility of placing the quantization module either before or after the Translator, we note that post-Translator placement induces significant training instability in classic VQ, attributable to the non-differentiable nature of the straight-through estimator. Conversely, our SVQ module demonstrates consistent stability during training. For consistency, the quantization module is placed before the Translator in our primary experiments, with a comparative assessment of its alternative placement provided in Section B.1.
4 Efficient Utilize of Cookbook Using Sparse Regression
To understand the difference between sparse regression-based quantization scheme and clustering-based quantization scheme, we measure the number of codes that we need in order to approximate any vector within an unit ball with error less than . We denote this number by . As indicated by the theorem below, using spare regression, can be reduced from to , with for high dimensional vectors.
@@ 4 Efficient Utilize of Cookbook Using Sparse Regression
for the clustering-based method is at least , while for sparse regression can be , with
as long as the non-zero elements used by sparse regression is at least
@@
To bound for the clustering method, we use the covering number for a unit ball , requiring at least code vectors to approximate any vector within error . Considering , with , , and an -sparse unit vector, we have:
implying
@@
with probability at least . Thus, . With the -sparse unit vector covering number bounded by , we obtain:
By choosing , we have
yielding , given .
As , it is easy to verify that , leading to . ∎
5 Experiments
We extensively evaluate SVQ on a wide range of real-world spatio-temporal forecasting datasets under the unified framework of OpenSTL (Tan et al., 2023b). Note that since SimVP holds leading performance across almost all benchmarks, it is adopted as our primary baseline. We begin by comparing the SimVP+SVQ model with state-of-the-art (SOTA) counterparts. Subsequently, we assess the capability of SVQ to enhance various backbones. Finally, we investigate the delicate balance between detail preservation and noise reduction, and conduct experiments with artificial noise injection to show the effect of SVQ on mitigating noise.
Moreover, we provide a comprehensive comparison between SVQ and existing VQ methods in Appendix B.3, evaluating their downstream prediction performance, codebook utilization, and computational efficiency.
@@ 5 Experiments
Dataset. To evaluate the proposed SVQ, we conduct extensive experiments on five real-world spatio-temporal forecasting tasks, including weather (WeatherBench), traffic flow (TaxiBJ), human pose dynamics (Human3.6M), driving scenes (KittiCaltech), and human actions (KTH Action). More details about datasets are described in Appendix A.1.
@@
Experimental details. In deployment, we found SVQ to be quite robust to codebook size, as its performance remains consistently strong when using a sufficiently large codebook. Therefore, we fix the codebook size at 10,000 for WeatherBench, TaxiBJ, and Human3.6M datasets, and at 6,000 for KittiCaltech and KTH datasets. The hidden dimension of nonlinear projection layer is fixed at 128. Our experiments are conducted on either 1 or 4 NVIDIA V100 32GB GPUs, with a total batch size of 16. More details about backbone architectures, VQ parameters, computational costs, and metrics are described in Appendix A.2, A.3, B.2, and A.5 respectively.
5.1 Benchmarks on various tasks
@@
| SimVP+SVQ (Frozen) | 1.023 | 0.6131 | 30.863 | 3.661 | 1.4337 | 0.7861 | 0.0456 | 0.1456 | |
| SimVP+SVQ (Learnable) | 1.018 | 0.6109 | 30.611 | 3.657 | 1.4186 | 0.7858 | 0.0458 | 0.1463 | |
@@
| SimVP+SVQ (Frozen) | 4.427 | 1.4160 | 360.15 | 12.445 | 23.915 | 3.4078 | 24.968 | 3.4117 | |
| SimVP+SVQ (Learnable) | 4.433 | 1.4164 | 360.53 | 12.449 | 23.908 | 3.4060 | 24.983 | 3.4095 | |

### v4

# Does Vector Quantization Fail in Spatio-Temporal Forecasting? Exploring a Differentiable Sparse Soft-Vector Quantization Approach

@@
Abstract.
Spatio-temporal forecasting is crucial in various fields and requires a careful balance between identifying subtle patterns and filtering out noise. Vector quantization (VQ) appears well-suited for this purpose, as it quantizes input vectors into a set of codebook vectors or patterns. Although VQ has shown promise in various computer vision tasks, it surprisingly falls short in enhancing the accuracy of spatio-temporal forecasting. We attribute this to two main issues: inaccurate optimization due to non-differentiability and limited representation power in hard-VQ. To tackle these challenges, we introduce Differentiable Sparse Soft-Vector Quantization (SVQ), the first VQ method to enhance spatio-temporal forecasting. SVQ balances detail preservation with noise reduction, offering full differentiability and a solid foundation in sparse regression. Our approach employs a two-layer MLP and an extensive codebook to streamline the sparse regression process, significantly cutting computational costs while simplifying training and improving performance. Empirical studies on five spatio-temporal benchmark datasets show SVQ achieves state-of-the-art results, including a 7.9% improvement on the WeatherBench-S temperature dataset and an average mean absolute error reduction of 9.4% in video prediction benchmarks (Human3.6M, KTH, and KittiCaltech), along with a 17.3% enhancement in image quality (LPIPS). Code is publicly available at https://github.com/Pachark/SVQ-Forecasting.
1. Introduction
Spatio-temporal forecasting is pivotal in numerous domains ranging from environmental monitoring to urban planning, where precisely predicting future dynamics is crucial. Among various methodologies explored in deep learning, Vector Quantization (VQ) has distinguished itself primarily in computer vision tasks, showcasing its ability to compress high-dimensional vectors into a compact, discrete form that maintains significant fidelity to the original information. Historically rooted in signal processing, VQ’s breakthrough came with its application in image processing advancements like the Vector Quantized-Variational AutoEncoder (VQ-VAE) (van den Oord et al., 2017), which set a precedent in generating high-quality images by learning efficient representations of complex distributions.
While VQ has proven effective and has become a nearly default approach in generation tasks, its potential in spatio-temporal forecasting remains underexplored. In spatio-temporal forecasting, data are considered including patterns, details, and noise. VQ offers the capability to capture recurring patterns through a discrete codebook, while simultaneously filtering out noise and irrelevant details that are unlikely to recur. Given its noise reduction properties and the similarities between image/video generation and spatio-temporal forecasting, one might reasonably expect VQ to yield performance gains in the latter. However, our review of existing studies reveals that few have successfully leveraged VQ to enhance forecasting accuracy. Our empirical evaluation of recent VQ methods shows that all fell short of expectations, often degrading the performance of baseline forecasting models rather than providing the expected improvements, as illustrated in Figure 1 and detailed in Table 5. They consistently exhibited a negative impact on final forecasting metrics such as mean squared error (MSE).
We believe that the similarity between image/video generation and spatio-temporal forecasting underscores the potential of VQ, but the problem lies in the inherent dynamic nature of spatio-temporal data. The complex temporal evolutions and spatial distributions introduce a level of complexity that traditional VQ methods are not equipped to handle. Specifically, the unsatisfactory results of traditional VQ methods are caused by two reasons:
Inaccurate model optimization caused by non-differentiability. The discrete nature of the quantization step prevents gradients from being directly passed through this operation. VQ methods typically employ the straight-through (or stop-gradient) estimator, as described in VQVAE (van den Oord et al., 2017). This estimator approximates the gradient by copying gradients from the quantized outputs to the input vectors, introducing errors in the optimization process.
Limited representation power of hard-VQ. The rapid evolution of spatiotemporal data over time, makes learning transformations from historical inputs to future outputs challenging. Hard-VQ struggles to replace the latent vector with the nearest codebook vector as the latent changes quickly, which limits the modeling of detailed spatio-temporal dynamics required for forecasting.
In response to these limitations, this work presents Differentiable Sparse Soft-Vector Quantization (SVQ), designed to strike a balance between noise reduction and detail preservation for spatio-temporal forecasting tasks. We solve the aforementioned challenges by:
Introduced a differentiable VQ mechanism to enable efficient gradient propagation. This involves using a two-layer MLP to approximate sparse regression, maintaining differentiability and enhancing accuracy. By generating regression coefficients through nonlinear projections and deriving quantized outputs using a codebook matrix, the approach facilitates gradient flow from outputs to inputs, addressing computational challenges in traditional VQ.
Developed Soft-VQ, a method that integrates sparse regression with a large codebook, allowing multiple vector allocations for each input. Compared to hard-VQ, SVQ effectively captures complex patterns while enhancing noise filtration. It preserves diverse information from inputs, better capturing dynamic features in spatio-temporal data. Remarkably, SVQ can utilize a static, randomly initialized codebook without performance degradation, thus reducing learning parameters and demonstrating efficiency and robustness.
Through extensive evaluation on a variety of real-world datasets, SVQ emerges as the first VQ method to achieve significant enhancements in spatio-temporal forecasting tasks. Notably, SVQ surpassed the leading model in WeatherBench-S temperature forecasting benchmark, achieving a 7.9% reduction in MSE. In video prediction tasks-Human3.6M, KTH, and KittiCaltech, SVQ systematically reduced the Mean Absolute Error (MAE) by 9.4%, and significantly improved perceptual quality, as indicated by a 17.3% reduction in the learned perceptual image patch similarity (LPIPS) score. These results underscore SVQ’s remarkable capability across a wide range of spatio-temporal forecasting scenarios.
2. Related Work
Here, we present a brief overview of vector quantization, the lineage of sparse coding techniques, and the latest developments in spatio-temporal forecasting models.
@@ 2. Related Work
Vector Quantization and Sparse Coding. Instead of using continuous latent, VQ-VAE (van den Oord et al., 2017), a seminal work, incorporates vector quantization to learn discrete latent representations, typically assigning each vector to the nearest code in a codebook. Subsequent enhancements include Residual VQ (Zeghidour et al., 2022), which quantizes the residuals recursively, and Multi-headed VQ (Mama et al., 2021a), which adopts multiple heads for each vector. While these methods are effective, they often rely on a relatively small number of codes to represent the original vectors. To address this, SCVAE (Xiao et al., 2023) employs sparse coding, allowing vectors to be represented through sparse linear combinations of multiple codes, and achieves end-to-end training via the Learnable Iterative Shrinkage Thresholding Algorithm (LISTA) (Gregor and LeCun, 2010). However, a significant drawback of the sparse coding method using LISTA (referred to as SVQ-raw here) is its high computational complexity, which scales quadratically with codebook size.
@@
Building on these insights and limitations, our work proposes a soft-VQ method applicable to spatio-temporal forecasting tasks. Although it is closely related to a simultaneous research work (Tschannen et al., 2023), which employs an infinite cookbook with a linear layer for continuous vector quantization applied in image generation, our approach is largely inspired by sparse regression, as clearly evidenced by our analysis. Specifically, our work focuses on the challenges arising from spatio-temporal forecasting, providing a strong theoretical foundation and effectively addressing the challenges.
Spatio-Temporal Forecasting Models. Recent advancements in spatio-temporal forecasting have highlighted a shift from recurrent to non-recurrent frameworks. Despite the forecasting capabilities of models like ConvLSTM (SHI et al., 2015), PredRNN (Wang et al., 2017), and PredRNNV2(Wang et al., 2022), this shift is largely due to the high computational demands of sequential processing in recurrent models. Non-recurrent models, such as MMVP (Zhong et al., 2023) and the SimVP family (Gao et al., 2022; Tan et al., 2022), have become benchmarks in video prediction by decoupling spatial and temporal learning through an efficient encoder-translator-decoder structure. This transition is further enhanced by innovative features like visual attention in TAU (Tan et al., 2023a) and MetaFormers in OpenSTL (Tan et al., 2023b), showcasing the continuous improvements towards more effective forecasting solutions. Our proposed method is designed for seamless integration as a plugin with the majority of these spatio-temporal forecasting models.
3. Differentiable Sparse Soft-Vector Quantization (SVQ)
We first present the mathematical foundation of differentiable sparse soft-vector quantization. Our proposed method effectively solves the optimization problem in a differentiable manner, and the theoretical analysis of codebook utilization demonstrates its strong representational capacity. We then describe its detailed implementation within a spatio-temporal forecasting model.
3.1. Vector Quantization by Sparse Regression
@@
where is the weight for combining codes to approximate . refers to the regularization parameter. By introducing regularizer in the optimization problem, we effectively enforce to be associated with a small number of codes. Compared to classic VQ methods where codes have to be learned through clustering, according to (Chiu et al., 2022), it is sufficient to use randomly sampled vectors as codes as long as its number is large enough, thus avoiding the need of computing and adjusting codes. The theoretical guarantee of sparse regression is closely related to the property of subspace clustering, as revealed in Theorem 3.1.
@@
where , is the th element of vector , sgn denotes the sign function, and outputs if and otherwise. We consider the first step of the iteration where and have and the resulting output for is given as
@@
3.2. Efficient Utilization of Cookbook Using Sparse Regression
To understand the difference between sparse regression-based quantization scheme and clustering-based quantization scheme, we measure the number of codes required to approximate any vector within a unit ball with error less than . This number is denoted by . Intriguingly, as the theorem below reveals, using sparse regression allows to be significantly reduced from to , where for high-dimensional vectors.
@@ 3.2. Efficient Utilization of Cookbook Using Sparse Regression
For the clustering-based method, is at least . In contrast, for sparse regression, can be formulated as , where
given that the number of non-zero elements utilized by sparse regression is at least
@@
To estimate for the clustering method, we consider the covering number for a unit ball which necessitates at least code vectors to approximate any vector within an acceptable error margin of . With where , and an -sparse unit vector, we discern:
which implies
@@
with probability at least . Therefore, , where denotes a covering vector. Since the -sparse unit vector covering number is bounded by , we establish:
Setting yields
therefore, , where . As long as , it follows that , affirming that . ∎
3.3. Spatio-Temporal Forecasting Model Enhanced by Quantization
Architecture of backbone model. As shown in Figure 3, SimVP (Tan et al., 2022) is employed as the backbone model, which encompasses an encoder for spatial feature extraction, a translator for temporal dependency learning, and a decoder for frame reconstruction. The quantization module is integrated between the encoder and translator. The input data is a 4D tensor , representing height (), width (), time step (), and channel (). The encoder condenses into downsampled latent representation , maintaining temporal dimensionality while altering spatial and channel dimensions. This latent space, composed of tokens, each represented by a -dimensional vector, undergoes vector quantization.
Quantization module. The SVQ comprises a two-layer MLP and an extensive codebook. The codebook is a randomly initialized matrix , where denotes the size of codebook. To achieve automatic selection of codes, a weight matrix is generated via nonlinear projection from the latent representation . This projection is formally expressed as , wherein the comprises two linear layers and an intermediate ReLU activation function. The quantized output is then obtained by computing the matrix multiplication of weight matrix and codebook matrix , a process that can be conceptualized as a selection operation as shown in Figure 3. To encourage sparsity within the generated weight matrix, we apply a Mean Absolute Error (MAE) loss to the output as a surrogate form of regularization.
4. Experiments
We evaluate SVQ under the unified framework of OpenSTL (Tan et al., 2023b).
@@ 4. Experiments
Datasets. We conduct experiments on five real-world spatio-temporal forecasting tasks, including weather (WeatherBench (Rasp et al., 2020)), traffic flow (TaxiBJ (Zhang et al., 2017)), human pose dynamics (Human3.6M (Ionescu et al., 2014)), driving scenes (KittiCaltech (Geiger et al., 2013; Dollár et al., 2009)), and human actions (KTH Action (Schüldt et al., 2004)). Details about datasets are provided in Appendix A.1.
@@
Experimental details. We found SVQ to be quite robust to codebook size, as its performance remains consistently strong when using a sufficiently large codebook. Therefore, we fix the codebook size at 10,000 for WeatherBench, TaxiBJ, and Human3.6M, and at 6,000 for KittiCaltech and KTH. The hidden dimension of nonlinear projection layer is fixed at 128. Experiments are conducted on 1 or 4 NVIDIA V100 32GB GPUs, with a total batch size of 16, and early stopped with a patience of 10. Details about backbone architectures and computational costs are described in Appendix A.2 and B.1.
Metrics. Forecasting accuracy is evaluated using mean squared error (MSE) and mean absolute error (MAE), while the image quality of predicted frames is assessed using structural similarity index measure (SSIM) (Wang et al., 2004), peak signal-to-noise ratio (PSNR), and learned perceptual image patch similarity (LPIPS) (Zhang et al., 2018).
Baselines. The comparison baselines include: 1) Non-recurrent models such as SimVP (Tan et al., 2022) and TAU (Tan et al., 2023a); 2) Recurrent-based models such as ConvLSTM (SHI et al., 2015), PredNet (Lotter et al., 2017), PredRNN (Wang et al., 2017), MIM (Wang et al., 2019b), E3D-LSTM (Wang et al., 2019a), PhyDNet (Guen and Thome, 2020), MAU (Chang et al., 2021), PredRNNv2 (Wang et al., 2022), and DMVFN (Hu et al., 2023). Baseline results are copied from OpenSTL (Tan et al., 2023b). For SimVP, we report the best MetaFormer variant on each dataset, detailed in Appendix A.2. It is worth noting that while some diffusion-based methods (Feng et al., 2023; Chen et al., 2023; Gupta et al., 2024) have shown strong performance in video generation, our focus is on spatiotemporal forecasting, which is typically evaluated using MSE and MAE rather than image authenticity. Therefore, we did not include them in our baselines. Further exploration is needed to understand their potential in forecasting settings, given the complex noise-signal interactions.
@@
| SimVP+SVQ (Frozen codebook) | 1.023 | 0.6131 | 30.863 | 3.661 | 1.4337 | 0.7861 | 0.0456 | 0.1456 | |
| SimVP+SVQ (Learnable codebook) | 1.018 | 0.6109 | 30.611 | 3.657 | 1.4186 | 0.7858 | 0.0458 | 0.1463 | |
@@
| SimVP+SVQ (Frozen codebook) | 4.427 | 1.4160 | 360.15 | 12.445 | 23.915 | 3.4078 | 24.968 | 3.4117 | |
| SimVP+SVQ (Learnable codebook) | 4.433 | 1.4164 | 360.53 | 12.449 | 23.908 | 3.4060 | 24.983 | 3.4095 | |
@@
4.1. Benchmarks on various forecasting tasks
| Model | MSE | MAE | SSIM | PSNR |
| ConvLSTM(SHI et al., 2015) | 0.3358 | 15.32 | 0.9836 | 39.45 |
| E3D-LSTM(Wang et al., 2019a) | 0.3427 | 14.98 | 0.9842 | 39.64 |
| PhyDNet(Guen and Thome, 2020) | 0.3622 | 15.53 | 0.9828 | 39.46 |
| PredNet(Lotter et al., 2017) | 0.3516 | 15.91 | 0.9828 | 39.29 |
| PredRNN(Wang et al., 2017) | 0.3194 | 15.31 | 0.9838 | 39.51 |
| MIM(Wang et al., 2019b) | 0.3110 | 14.96 | 0.9847 | 39.65 |
| MAU(Chang et al., 2021) | 0.3268 | 15.26 | 0.9834 | 39.52 |
| DMVFN(Hu et al., 2023) | 3.3954 | 45.52 | 0.8321 | 31.14 |
| PredRNN++(Wang et al., 2018) | 0.3348 | 15.37 | 0.9834 | 39.47 |
| PredRNN.V2(Wang et al., 2022) | 0.3834 | 15.55 | 0.9826 | 39.49 |
| TAU(Tan et al., 2023a) | 0.3108 | 14.93 | 0.9848 | 39.74 |
| SimVP (w/o VQ)(Tan et al., 2022) | 0.3246 | 15.03 | 0.9844 | 39.71 |
| SimVP+SVQ (Frozen codebook) | 0.3171 | 14.68 | 0.9848 | 39.83 |
| SimVP+SVQ (Learnable codebook) | 0.3191 | 14.64 | 0.9849 | 39.86 |
| Improvement | 1.7% | 2.6% | 0.1% | 0.4% |